OpenStudy (anonymous):

a university student sells 1600 cups of coffee per day at a price of $2.40. a market survey shows that for every $0.05 reduction in price, 50 more cups of coffee will be sold. how much should the student center charge for a cup of coffee in order to maximize revenue?

OpenStudy (anonymous):

Revenue is quantity times price.

OpenStudy (anonymous):

R(x)=x*p(x)

OpenStudy (anonymous):

x = price reductions

OpenStudy (anonymous):

Let x be the number of (5c) discounts. Lets use cents as units for revenue. Then: Revenue = R = Volume x price = (1600+50x)(240-5x) R = 384000 +4000 x -250 x^2 To maximise revenue dR/dx = 0 dR/dx = 4000-500x=0 x=8 I.e. 8 discounts of 0.05 cents

OpenStudy (anonymous):

@iambatman then \(x\) isn't quantity?

OpenStudy (anonymous):

Which means 2.40 - 8x0.05 = $ 2 for a cuppa

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

what is \(xp(x)\) then?

OpenStudy (anonymous):

x = number of price reductions the price of a cup of coffee -> p =2.4-0.05x Number of cups of coffee sold -> 1600+50x Revenue = R(x) = x*p(x) = (1600+50x)(2.4-0.05) = -2.5x^2+40x+3840 R'(x) = 40-5x = 0 5x = 40 x = 8 p = 2.40-0.05(8) = $2.00 a cup